OpenStudy (anonymous):

A compound of C,H and N contains 66.70 of C 7.41% of H, 25.90% of N ,,the molecular mass of compound was found to be 108...find its empirical formula and molecular formula

OpenStudy (anonymous):

Question is almost done but just Vapor Density of the compound is required. Without it, the data is insufficient.

OpenStudy (anonymous):

First assume an original sample of 100 grams- \[66.70/100=.667g\] \[7.41/100=.0741g\] \[25.90/100= .259g\] Now convert grams to moles \[.667/12.01=.0555mol\] \[.0741/1.008=.0735mol\] \[.259/14.01=.0185mol\] Now divide each by the smallest amount \[.0555/.0185=3\] \[.0735/.0185=4\] \[.0185/.0185=1\] Use these numbers as the formula numbers for the empirical formula- \[C _{3}H _{4}N\] Now calculate the molar mass of this molecule 54.07g And divide the experimental mass by this one- \[108/54.07=2\] Now multiply each of the empirical formula units by this number- \[C _{6}H _{8}N _{2}\] And that's it.